Let f(x)=x^2+x and g(x)=2x+1. Find and simplify the following:
.
a) (g o f)(x)
.
b) (f * g)(x) -- the o is colored in.
.
c) (f+g)(2)
.
d) (f-g)(3)

Answer by glitzgirl_14(11)

f(x)=x^2 g(x)=2x+1
A.) g(f(x))= 2(x^2)+1 (g o f)(x)=2x^2+1
B.) x^2(2x+1) 2x^3+x^2
C.) x^2+2x+1
D.) x^2-2x+1